CourtneyWatson:

Myra owns a car service that charges a $5 flat rate and an additional $0.50 per mile, which is represented by the equation y = 0.5x + 5, where x is the number of miles and y is the total cost. How much is the total cost for a car with 30 miles? $10 $15 $20 $35

mxddi3:

ok so you see how it says X is number of miles? and the question says there are 30 miles? well put 30 in place of x and solve it. does this make sense or do you need a visual?

ramen:

Since we already have the equation, the problem states "an additional $0.50 per mile" so this means that we would plug in that 30 miles into x. So this makes y=.5(30)+5

CourtneyWatson:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mxddi3 ok so you see how it says X is number of miles? and the question says there are 30 miles? well put 30 in place of x and solve it. does this make sense or do you need a visual? \(\color{#0cbb34}{\text{End of Quote}}\) yes it makes sense

CourtneyWatson:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen Since we already have the equation, the problem states "an additional $0.50 per mile" so this means that we would plug in that 30 miles into x. So this makes y=.5(30)+5 \(\color{#0cbb34}{\text{End of Quote}}\) Ok

justus:

y = 0.5x +5 y = 0.5(30) + 5 y = 15 + 5 Just add 15 + 5 to get your answer.

CourtneyWatson:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @justus y = 0.5x +5 y = 0.5(30) + 5 y = 15 + 5 Just add 15 + 5 to get your answer. \(\color{#0cbb34}{\text{End of Quote}}\) Oh ok thank you so its $20

ramen:

Yes that is the correct answer good job.

justus:

Correct, you're most welcome.

justus:

Thank you for everyone else who was assisting the user. :)

ramen:

No problem.

justus:

You can now close the post if you'd like.

CourtneyWatson:

Thank you @mxddi3 @ramen and @justus for helping me

justus:

:)